OpenStudy (nerdgirl):

Find the sum of the following series. (shown below) A. 240 B. 255 C. 210 D. 510

OpenStudy (nerdgirl):



OpenStudy (nerdgirl):

@aloud that's not one of my answers @phi PLEASE HELP!

OpenStudy (nerdgirl):

no you ditz it's not one of them don't you get it?! it's not one of the answer choices!

OpenStudy (nerdgirl):

@phi

OpenStudy (phi):

rewrite the problem \[\sum_{1}^{15}(2n+1)= 2\sum_{1}^{15}n+\sum_{1}^{15}1\]

OpenStudy (phi):

There is a formula to add up the numbers from 1 to n do you know it? and the second sum means add up 15 1's (which hopefully you know how to do)

OpenStudy (nerdgirl):

Mmm.... what would you say if I said I knew it but I couldn't remember it? :p

OpenStudy (phi):

Gauss (famous mathematician), when a young kid in school was given the problem of adding up the numbers from 1 to 100, and he (clever fellow) saw a way to do it quickly. \[\sum_{k=1}^{n}k= \frac{ n(n+1) }{ 2 }\]

OpenStudy (phi):

your problem has n=15 so use that formula with n=15 and n+1 = 16

OpenStudy (nerdgirl):

so it would be 15 (greek symbol) k=1 (1) = 15 (16)/2? @phi

OpenStudy (nerdgirl):

@phi

OpenStudy (phi):

the sum of the numbers 1 to 15 is 15*16/2 = 15*8= 120 \[ \sum_{1}^{15}(2n+1)= 2\sum_{n=1}^{15}n+\sum_{n=1}^{15}1 \\ = 2(120) + \sum_{1}^{15}1 \] the sum of 15 ones is 15*1= 15 thus \[ \sum_{1}^{15}(2n+1)= 2\sum_{n=1}^{15}n+\sum_{n=1}^{15}1 \\ = 2(120) + 15 \] can you finish?

OpenStudy (nerdgirl):

OMG ILY THANK YOU!!!!!!!!!!!